Session: Nov 1997 (session 4)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Tue Nov 18 01:41:54 MET 1997) Success
No known problems

Onsala (A.Polatidis at Tue Nov 18 11:19:53 MET 1997) Success
No known problems. We are recording only LCP so expect signal on ODD VCs.

Noto (C.Trigilio at Tue Nov 18 13:43:02 MET 1997) Success
No problem. Cloudy sky 100%.

Effelsberg (dgraham at Fri Nov 21 09:42:29 MET 1997) Minor failures
rather damp. Tsys values scatter due to tcal too small.

Jodrell (Alastair Gunn at Sat Nov 22 19:41:03 MET 1997) Minor failures
Experiment started late due to reciever problems. Data good from pass 3. Formatter re-synched at about 18:00 but discovered to be 2 seconds behind GPS and reset at 18:23. Telescope taken offline at 19:35 due to high winds - no data after this time! Therefore tape IAA00157 contains no data.

Metsahovi (Kaj Wiik at Mon Nov 24 13:19:17 MET 1997) Success
WX cloudy. Spurious ERROR messages due to too long comment line in procedure file.

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:06:26 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name el22b.antab Read the comments in the file to get further information.